An RLC circuit is used in a radio to tune into an FM station broadcasting at f = 99.7 MHz. The resistance in the circuit is R = 12.6 Ω, and the inductance is L = 1.70 µH. What capacitance should be used?
1
Expert's answer
2018-09-26T10:22:08-0400
The resonance frequency of an RLC circuit
ω=2πf=1/√LC
Therefore
C=1/(4π^2 f^2 L)
=1/(4×〖3.14〗^2×(99.7×〖10〗^6 Hz)^2×1.70×〖10〗^(-6) H)=1.50×〖10〗^(-12) F=1.50 pF